# RotBench: Evaluating Multimodal Large Language Models on Identifying Image Rotation

### v2

@@ Appendix D Clockwise Prompting
Appendix E Differentiating Clockwise and Counter-clockwise Rotations using GPT-5
@@
In Table˜7, we repeat the binary clockwise versus counter-clockwise classification experiment (Section˜6.2) using GPT-5. GPT-5 achieves considerably higher accuracy on this binary classification task compared to GPT-4o and Qwen-2.5-VL-7B-Instruct. Notably, performance on 90° and 270° images are significantly higher in the binary task than in the four-way classification task.
@@
Appendix F Chain-of-Thought Example
@@
Appendix G Modifying Temperature
@@
To ensure model performance does not significantly vary with sampling temperature, we examine zero-shot performance of GPT-4o and Qwen-2.5-VL-7B-Instruct (Qwen) on varying temperature settings (Table˜8). We see that performance on all four rotations does not vary significantly for GPT-4o. Qwen does show slight sensitivity to temperature. As temperature increases, performance on 0° and 90° decreases, while performance on 180° and 270° increases. These results support our primary conclusions, suggesting that the model distinguishes between landscape rotations (0° and 180°) and portrait rotations (90° and 270°). The model defaults to 0° for landscape rotations, and 90° for portrait rotations. With temperature increases, the model’s responses begin to resemble random guessing.
Appendix H In-Context Learning
@@
Results.Table˜9 shows the accuracy on each image orientation after injecting in-context examples. We do not see consistent performance improvement, regardless of the number of in-context examples provided. Together with results in Table˜1, these experiments suggest that various forms of prompt modification are insufficient for identifying rotation. Rather, robust orientation identification demands explicit parameter optimization through fine-tuning.
Appendix I Normalized Rotation Voting
Accuracy (↑)
Model 90° 180° 270°
Qwen-2.5-VL-7B-Instruct
Zero-shot 0.95 0.57 $0.03$ $0.15$
Chain-of-Thought $0.86$ $0.20$ $0.13$ $0.30$
Normalized Rotation Voting $0.54$ $0.52$ 0.56 0.52
GPT-4o
Zero-shot $0.87$ $0.65$ $0.85$ $0.21$
Chain-of-Thought 0.91 $0.59$ 0.86 $0.21$
Normalized Rotation Voting $0.86$ 0.88 $0.80$ 0.86
Algorithm 1 Normalized Rotation Voting
[…]
3:for $i\leftarrow 0$ to |images|1|images|-1 do
4:  $rot\leftarrow\text{call\_model}(images[i])$
5:  $rot\_norm\leftarrow(rot-i\times 90)\bmod 360$
6:  append rot_normrot\_norm to rotationsrotations
7:end for
[…]
Intuition and Setup.  Our results show frontier MLLMs are able to reliably identify 0° and 180° images. Leveraging this pattern, we propose a voting approach to identify image rotation. Given an input image of unknown orientation, we can further rotate this image 0°, 90°, 180°, and 270°. We separately prompt the model to identify the rotation of each image, then normalize the predictions by subtracting the applied rotation (effectively shifting the angle into a common reference frame) (Algorithm˜1). Regardless of the original rotation, two of the further rotated images would correspond to 0° and 180° in the ground-truth reference frame. If the model can correctly identify these two images, there is a high probability that the majority vote would reveal the ground-truth rotation. We evaluate GPT-4o and Qwen-2.5-VL-7B-Instruct on RotBench-small using this approach.
Results.Table˜10 shows the accuracy on each image orientation using normalized rotation voting. In general, we see a much more even performance distribution across all four rotations compared to using zero-shot or chain-of-thought prompting. Qwen-2.5-VL-7B-Instruct achieves substantially stronger performance on 0° and 90° (0.95 and 0.57) compared to 180° and 270° (0.03 and 0.15). Normalized rotation voting achieves $0.5$ performance on all orientations. GPT-4o achieves $0.8$ accuracy on all orientations. In particular, performance on 270° improved substantially compared to zero-shot prompting (+0.85+0.85).
Appendix J Similar Images in Spatial-MM
@@ Appendix I Normalized Rotation Voting
During the data filtering process, we noted several images in Spatial-MM closely resemble each other (Fig.˜11). This realization led to our decision to use MS COCO as the training dataset in our fine-tuning experiment.
Appendix K License
We will publicly release our code and models. We provide the following links to the standard licenses for the datasets, code, and models used in this project.
Appendix L Prompts
This section includes all prompts used in our experiments.
L.1 Extract Image Subjects
Fig.˜12 describes the prompt used for extracting primary subjects in images. The list of subjects extracted is later used to obtain bounding boxes, scene graphs, and segmentation maps .
@@
L.2 Captions
Fig.˜13 describes the prompt used for captioning images.
@@
L.3 Scene Graphs
Fig.˜14 describes the prompt used for extracting scene graphs from images.
@@
L.4 Clockwise Experiment
Fig.˜15 describes the prompt used for clockwise vs counter-clockwise rotation experiment.
System Prompt You are an intelligent AI assistant that specializes in identifying rotation in images. You will be given an image that has been rotated 90 or 270 degrees. Specifically, a 90 degree rotation is a quarter-turn counter-clockwise (the same as 270 degrees clockwise); 270 is three quarter-turns counter-clockwise (the same as 90 degrees clockwise).
@@
L.5 Rotation Classification
Fig.˜16 describes the system and user prompts for rotation classification. The mapping between letter choice and degrees is shuffled each prompt.

### v3

@@ Appendix F Clockwise Prompting
Appendix G Differentiating Clockwise and Counter-clockwise Rotations using GPT-5
@@
In Table˜9, we repeat the binary clockwise versus counter-clockwise classification experiment (Section˜6.2) using GPT-5. GPT-5 achieves considerably higher accuracy on this binary classification task compared to GPT-4o and Qwen-2.5-VL-7B-Instruct. Notably, performance on 90° and 270° images are significantly higher in the binary task than in the four-way classification task.
@@
Appendix H Chain-of-Thought Example
@@
Appendix I Modifying Temperature
@@
To ensure model performance does not significantly vary with sampling temperature, we examine zero-shot performance of GPT-4o and Qwen-2.5-VL-7B-Instruct (Qwen) on varying temperature settings (Table˜10). We see that performance on all four rotations does not vary significantly for GPT-4o. Qwen does show slight sensitivity to temperature. As temperature increases, performance on 0° and 90° decreases, while performance on 180° and 270° increases. These results support our primary conclusions, suggesting that the model distinguishes between landscape rotations (0° and 180°) and portrait rotations (90° and 270°). The model defaults to 0° for landscape rotations, and 90° for portrait rotations. With temperature increases, the model’s responses begin to resemble random guessing.
Appendix J The Limitations of Scene Graphs
90° 180° 270°
- 0.33 0.16 0.36
90° 0.33 - 0.21 0.20
180° 0.16 0.21 - 0.20
270° 0.36 0.20 0.20 -
Table 11: Rate of overlapping scene graphs extracted from images rotated different orientations.
Our primary results (Section˜5) show that including various forms of auxiliary information fails to consistently improve model performance. Our intuition behind providing scene graphs is that it would capture unnatural spatial relationships between objects (i.e., “horse above human”). However, strong language priors may override the image understanding, resulting in the model providing the same caption or scene graph regardless of image rotation. To understand why scene graphs fail to improve performance, we examine the overlap between scene graphs extracted from images rotated different orientations. We define two scene graphs as overlapping when both subjects and the predicate are identical, and when both the order of subjects are reversed and the spatial predicate is inverted (e.g., ["cow", "left", "farmer"] and ["farmer", "right", "cow"] are identical scene graphs).
Table˜11 displays the rate of overlapping scene graphs between different image orientations. Noticeably, there is a much higher degree of overlap between 0° and 90°/270° (around 0.3) compared to between 0° and 180° (0.16). These results show the caption model is able to recognize upside-down image and caption accordingly, but fails to correctly capture vertically rotated images.
Appendix K Alternative Evaluation Methods
Model Name MAAE ($\downarrow$)
Qwen 2.5 VL 7B Instruct $88.72\pm 2.53$
GPT-4o $61.09\pm 0.48$
Table 12: Comparison of Mean Absolute Angular Error (MAAE) for Qwen-2.5-VL-7B-Instruct and GPT-4o on a regression-based rotation identification task.
Our primary experiments (Table˜1, Table˜2) both follow the four-option multiple-choice setup widely used in previous literature (Anis et al., 2025; Joshi and Guerzhoy, 2017; Udandarao et al., 2024) to enable a more direct comparison and analysis with previous results. In this section, we explore model performance on alternative evaluation schemes. We see that all alternative setups show inferior performance. We select our current 4-option multi-choice setup for our primary experiments as it shows that, despite the leniency of the setup, frontier models still perform unsatisfactorily.
Regression.  Given that the current setup already challenges frontier reasoning models, we expect evaluating the models on a regression task on the continuous interval $0^{\circ},359^{\circ}]$ will result in further degraded performance. In fact, doing so may actually limit the conclusions and insights we can draw from our experiments as it would be difficult to infer any failure cases without first discretizing the interval. To illustrate this, we provide some further experiments evaluated Qwen-2.5-VL-7B-Instruct and GPT-4o using a modified regression setup. Each image in RotBench-large is randomly rotated $0^{\circ}-359^{\circ}$ counter-clockwise. The model is asked to output a single integer representing the degree of rotation. We then calculate the Mean Absolute Angular Error (MAAE) between the estimated and true rotations. Compared to an absolute error, MAAE takes into account the cyclic nature of angular measurements:
$\mathrm{MAAE}=\frac{1}{n}\sum_{i}^{n}\min\!\left(|\hat{\theta}_{i}-\theta_{i}|,\,360-|\hat{\theta}_{i}-\theta_{i}|\right)$
We find that both models are unable to provide accurate predictions of image orientation. GPT-4o performed significantly better than Qwen-2.5-VL-7B-Instruct, with a difference in MAAE of about 20. Nonetheless, both models averaged >45>45 MAAE across the dataset, indicating a limited ability of current MLLMs to reason over continuous angular spaces.
Number of Bins Accuracy
Qwen-2.5-VL-7B-Instruct
4 $0.26\pm 0.03$
8 $0.16\pm 0.01$
10 $0.15\pm 0.01$
18 $0.10\pm 0.01$
GPT-4o
4 $0.38\pm 0.02$
8 $0.28\pm 0.01$
10 $0.23\pm 0.02$
18 $0.17\pm 0.02$
Table 13: Classification accuracy on RotBench-large using different discretization levels. Rows represent the number of bins, which is equivalent to the number of choices in the multiple-choice question.
Discretization at various granularities.  Alternatively, we evaluate how classification accuracy changes when discretizing the continuous interval $0^{\circ},359^{\circ}]$ into various granularities. For each image in RotBench-large, we randomly sample a rotation angle in [0°, 359°]. We then discretize the interval into a certain number of bins, assign a letter choice to each bin, and map the ground truth rotation angle to the correct letter choice. The models are then prompted to answer a multiple-choice question, where the number of choices equals the number of bins, and each choice represents a range of angles (e.g., “A. 0°-20°”). Doing so transformed our previous experiment from a regression task into a classification task. Note that our original experimental setup is identical to the four-bin setup.
We experiment with 4 bins (each representing 90°), 8 bins (45°), 10 bins (36°), and 18 bins (20°). As expected, we see accuracy worsen as we increase the number of bins (Table˜13). These results suggest that both evaluated models still lack the ability to reason about fine-grained image orientations.
Appendix L In-Context Learning
@@
Results.Table˜14 shows the accuracy on each image orientation after injecting in-context examples. We do not see consistent performance improvement, regardless of the number of in-context examples provided. Together with results in Table˜1, these experiments suggest that various forms of prompt modification are insufficient for identifying rotation. Rather, robust orientation identification demands explicit parameter optimization through fine-tuning.
Appendix M Similar Images in Spatial-MM
@@
During the data filtering process, we noted several images in Spatial-MM closely resemble each other (Fig.˜11). This realization led us to use MS COCO—an out-of-distribution dataset—as the training dataset in our fine-tuning experiment.
Appendix N Qualitative Examples of Easy and Difficult Images
Figure 12: Qualitative examples of images that are difficult for humans (left), difficult for GPT-5 and Gemini-2.5-Pro (center), and easy for both models (right).
This section provides a qualitative illustration of classification difficulty by presenting images that were misclassified by human annotators during Stage 2, alongside examples that both GPT-5 and Gemini-2.5-Pro classified correctly and incorrectly.
Human errors.  On RotBench-small, humans achieve 0.99 accuracy on images rotated 0°, 90°, 180°, and 0.97 accuracy on images rotated 270°. These results indicate that for each rotation, all the annotators combined made fewer than 2 misclassifications. The images that led to misclassifications typically display subjects that can be logically oriented in multiple orientations and require incorporating background information to make an accurate judgment, such as the bottom-left image in Fig.˜12.
Model errors.  To better understand how the models perceive difficulty, we examined the set of images GPT-5 and Gemini-2.5-Pro answered correctly when rotated 90°, 180°, and 270° (easy images), and the images they answered incorrectly in all rotations (hard images). We include two example images from each category in the center and right columns of Fig.˜12.
We find only 4 images in RotBench-small (50 images total) elicit incorrect classifications for both GPT-5 and Gemini-2.5-Pro across the three rotations. The images tend to have these characteristics:
  1. 1.
    The image features a prominent primary subject and blurred/subtle background. Moreover, the primary subject can be viable even when rotated in multiple orientations (bottom-center image in Fig.˜12). Accurate identification of rotation requires incorporating the background into the reasoning.
  2. 2.
    The rotational cue is slight or subtle. The ‘Flagged Image’ example in Fig.˜7 belongs to this category, as the sole rotational cue is the slight tilt of the camera. If the camera was directly looking down, the image would be viable in any rotation.
On the other hand, there are 5 images that both models got correct in all orientations. These images often feature a clear central character, or tend to be natural landscape images featuring clear color distinctions between sky and ground (right column in Fig.˜12).
Appendix O Licenses
We will publicly release our models, and include our code and data in the supplementary. We provide the following links to the standard licenses for the datasets, code, and models used in this project.
Appendix P Prompts
Fig.˜13 describes the prompt used for extracting primary subjects in images. The list of subjects extracted is later used to obtain bounding boxes, scene graphs, and segmentation maps. Fig.˜14 describes the prompt used for captioning images. Fig.˜15 describes the prompt used for extracting scene graphs from images. Fig.˜16 describes the prompt used for clockwise vs counter-clockwise rotation experiment. Fig.˜17 describes the system and user prompts for rotation classification. The mapping between letter choice and degrees is shuffled each prompt.
@@
System Prompt You are an intelligent AI assistant that specializes in identifying rotation in images. You will be given an image that has been rotated 90 or 270 degrees. Specifically, a 90° rotation is a quarter-turn counter-clockwise (the same as 270 degrees clockwise); 270 is three quarter-turns counter-clockwise (the same as 90° clockwise).